On planet number 3 you fire a projectile horizontally from the edge of a vertical cliff 80m above the ground with an initial speed of 60.2 m/s. The projectile lands 5.81 m away from the base of the cliff. What is the value of g for the planet number 3
Answer #1

Solution) vertical cliff height Y = 80 m

Initial horizontal velocity Vx = 60.2 m/s

Horizontal distance X = 5.81 m

g value for planet number 3 = ?

We have

Y = Vy(T) + (1/2)gT^2

Here Vy is initial vertical velocity and Vy = 0 as there is no initial vertical velocity since projectile is fired in horizontal direction.

T is time

g is acceleration of planet number 3

- 80 = 0 + (1/2)(-g)T^2

T^2 = (80×2)/g

T = (160/g)^(1/2)

We have

Horizontal displacement

X = Vx(T) + (1/2)ax(T^2)

Here ax is horizontal acceleration =0

X = Vx(T)

X = Vx(160/g)^(1/2)

5.81 = 60.2(160/g)^(1/2)

(g/160)^(1/2) = (60.2/5.81)

Squaring on both sides

(g/160) = (60.2/5.81)^2

(g/160) = 107.35

g = 160 × 107.35 = 17176 m/s^2